OpenStudy (adll23):

Solve 4(x + 1) = 4x + 10. (4 points) 0 6 All real numbers No solution

OpenStudy (alexandervonhumboldt2):

first distribute the 4*(x+1). what is 4*(x+1) @adll23 ?

OpenStudy (adll23):

all I know is it turns into 4x+4 I think

OpenStudy (alexandervonhumboldt2):

yes you are correct now we have 4x+4=4x+10 subtarct 4x from both sides: 4x-4x+4=4x+10-4x 0+4=10 4=10 which is false thus .........?

OpenStudy (adll23):

there is no solution

OpenStudy (alexandervonhumboldt2):

yes you are correct no solution

OpenStudy (adll23):

thank you!

OpenStudy (alexandervonhumboldt2):

np :)